# RNNs, CNNs and Transformers in Human Action Recognition: A Survey and A Hybrid Model

### v1

@@ 3.1 CNN-based approaches in HAR
In 2014, Simonyan et al. [35] presented an innovative approach to recognizing actions in video sequences by using a two-stream CNN architecture. This approach divides the task into two distinct problems: recognizing spatial features from single frames and capturing temporal features across frames. The spatial stream CNN processes static visual information, while the temporal stream CNN handles motion by analysing optical flow. The model was tested on benchmark datasets like UCF-101 and HMDB-51, where it achieved state-of-the-art results, showcasing the effectiveness of this dual-stream method. The novelty of this work lies in the separation of motion and appearance features, which allows for more specialized networks that can better capture the complexities of video-based action recognition. The success of this model has made a significant impact on the field, influencing many future research directions in video understanding.
Sun et al. [56] presented a novel approach to action recognition by integrating deep learning with Slow Feature Analysis (SFA). SFA is a computational technique used to extract features from a signal that change slowly over time. It is based on the principle that while the sensory data may change rapidly, the underlying causes of the data often change at a slower rate. By focusing on these slowly varying features, SFA aims to capture the inherent properties of the observed data that are more stable and consistent over time. In the context of visual data, such as video sequences, SFA would aim to identify features that remain constant or change slowly as the video progresses. These could be the overall shape of an object, its identity, or its trajectory through space, as opposed to the rapidly changing texture or lighting conditions. The authors developed a CNN that processes video data to extract a hierarchy of complex features for action recognition tasks. The research stands out for applying a deep learning structure to SFA, which traditionally uses a single-layer approach. This deepened SFA is shown to be superior in learning more generalizable features from video sequences. The model was evaluated on Hollywood2, KTH, and UCF Sports datasets, which are standard benchmarks with considerable intra-class variation challenges. The paper reports that the deep learning-enhanced SFA model surpasses traditional methods and some state-of-the-art techniques in accuracy, without relying on additional processing tricks such as dense sampling. This suggests that the model’s performance is due to its robust feature learning capabilities.
Ijjina et al. [57] presented a novel approach to HAR by leveraging the patterns found in action bank features through a specially designed CNN. Action bank features are computed by considering a fixed set of videos known as an action bank. These videos serve as templates to measure the similarity of new videos against them. The CNN is uniquely tailored with specific masks to capture these patterns within the UCF50 dataset. The novelty lies in this approach of using action bank features for action recognition. The results show that the CNN effectively learns from the data, with a decreasing error rate over iterations, indicating good generalization. The paper suggests the potential for applying this method to other datasets and exploring different CNN models.
In 2015, the use of CNNs in action recognition tasks was at the forefront of research in computer vision, particularly due to their ability to learn hierarchical feature representations directly from raw data. The field was rapidly evolving, with new architectures and techniques being proposed to leverage the power of CNNs for this complex task.
Cheron et al. [58] presented an innovative approach to action recognition in videos. The core contribution is the development of the Pose-based CNN (P-CNN) descriptor, which uniquely integrates motion and appearance information along the trajectories of human body parts. This methodology marks a departure from traditional action recognition techniques that predominantly rely on statistical representations of local motion descriptors. By focusing on the detailed analysis of individual body parts through CNNs, the P-CNN model demonstrates a heightened sensitivity to subtle and fine-grained actions, a notable advancement in the field. The paper validates the efficacy of this approach using challenging datasets like JHMDB and MPII Cooking Activities, where it shows a marked improvement over existing methods, underscoring the potential of pose-based features in enhancing the accuracy of action recognition systems.
Geng et al. [59] addressed the limitations of traditional classifiers that rely on complex, handcrafted features, the authors propose a model that utilizes CNNs to directly process raw video inputs. A key innovation of their work is the introduction of a pre-training strategy employing a Convolutional Auto-Encoder (CAE), aimed at reducing the computational burden associated with kernel training. This method not only facilitates more efficient training but also enhances the model’s ability to generalize from raw data, circumventing the need for manually designed features. Evaluated on the KTH action database, the proposed model demonstrates superior performance in recognizing human actions compared to existing algorithms that depend on handcrafted features, showcasing its potential for real-world applications in video surveillance, virtual reality, and intelligent human-computer interactions. This approach marks a significant advancement in the field of computer vision and deep learning, offering a robust solution to the challenges of HAR.
Du et al. [60] propose a novel, CNN-based framework for action recognition, leveraging the dynamics of human skeletal movements. This research departs from traditional methods by transforming skeleton sequences into matrices, which are then converted into images, encapsulating both spatial structures and temporal dynamics. These images are fed into a specially designed CNN, where simple max-pooling plays a crucial role in spatial feature selection and temporal frequency adjustment. This approach not only preserves the intricate motion characteristics of each body part but also addresses the challenges posed by variable sequence lengths and action frequencies. The model’s efficacy is demonstrated through its superior performance on the Berkeley MHAD [61] and the ChaLearn gesture recognition [62] datasets, where it achieves state-of-the-art results, particularly excelling with a significant margin over existing methods in terms of computational efficiency and accuracy. This advancement in skeleton-based action recognition represents a significant leap forward, offering robust, efficient solutions for applications in human-computer interaction, surveillance, and beyond.
Gkioxari et al. [63] introduced a novel approach to action recognition in static images, significantly enhancing the accuracy by incorporating contextual cues from the surrounding environment. The core innovation lies in the adaptation of the Region-based Convolutional Network (RCNN) to the RCNN model, which uniquely utilizes both a primary region, encompassing the person performing the action, and a secondary region that automatically identifies relevant contextual information. This dual-region approach allows R*CNN to capture a more comprehensive understanding of the action by considering additional elements like objects, other people, and the scene itself. The model’s effectiveness is demonstrated through its remarkable performance on the PASCAL VOC Actions dataset, where it achieves a mean AP of 90.2%, significantly surpassing existing methods. Additionally, its versatility is showcased by its application to the MPII Human Pose dataset and the Berkeley Attributes of People dataset, where it achieves state-of-the-art results in attribute classification. This paper’s contribution is pivotal, as it not only advances the field of action recognition by integrating contextual understanding into the recognition process but also opens avenues for fine-grained classification tasks, underscoring the model’s adaptability and robustness.
The year 2016 witnessed a surge in the use of CNN models for action recognition tasks. Numerous models emerged, employing multiple streams and modalities. Additionally, there was a notable development of models utilizing 3D-CNNs, as well as hybrid models that combined 2D-CNNs with RNNs to analyze temporal information in videos.
Bilen et al. [64] introduced a groundbreaking approach to video analysis, particularly for action recognition, using CNNs. The central innovation is the concept of ”dynamic images,” a compact video representation that encodes the temporal evolution of video frames into a single RGB image. This is achieved through an efficient process known as rank pooling. The paper presents a novel static/dynamic neural network architecture that enables end-to-end training from videos. This architecture uniquely combines static appearance information from still frames with dynamic information from the entire video sequence. A significant aspect of this model is the introduction of an approximate rank pooling CNN layer, which greatly speeds up the process compared to traditional rank pooling. The paper claims state-of-the-art performance in action recognition on standard benchmarks (HMDB51 and UCF101). The dynamic images primarily focus on the acting objects in the video, effectively summarizing actions and motions. However, the compression of video content into a single image, while efficient, might oversimplify complex actions or subtle nuances in videos.
Zhang et al. [65] introduces a significant advancement in the field of video-based action recognition by optimizing the deep two-stream architecture for real-time processing. The core innovation lies in replacing the computationally intensive optical flow with motion vectors, which are readily available from compressed videos, thus eliminating the need for additional calculations. However, motion vectors inherently lack the fine-grained structural details and contain noisy motion patterns, which could degrade recognition performance. To address this, the authors propose a novel approach of transferring knowledge from an optical flow CNN (OF-CNN) to a motion vector CNN (MV-CNN). This transfer is achieved through three strategies: initialization transfer, supervision transfer, and a combination of both. By doing so, the enhanced MV-CNN effectively learns from the more accurate motion patterns captured by OF-CNN, leading to a significant boost in performance. The paper demonstrates that this method achieves comparable recognition accuracy to state-of-the-art techniques while processing at an impressive speed of 390.7 frames per second, which is 27 times faster than the original two-stream method. This breakthrough not only provides a feasible solution for real-time action recognition but also opens up new possibilities for efficient video analysis in various applications. The evaluation of this method on large datasets like UCF101, HMDB51, and THUMOS14 further underscores its effectiveness and practical utility in real-world scenarios.
Rahmani et al. [66] introduced a novel approach for recognizing human actions in 3D from various viewpoints, a significant challenge in video-based HAR due to factors like viewpoint variations, illumination, and visual appearance. Central to their methodology is a deep CNN that transforms human poses from different, unknown views into a view-invariant high-level space. Uniquely, the training data for this model is synthetically generated by fitting 3D human models to real motion capture data and rendering these poses from multiple viewpoints, circumventing the need for extensive and expensive multiview training data. This approach is innovative in its use of synthetic data for training while maintaining the ability to generalize to real depth images of unseen poses without re-training. The paper also introduces a group sparse Fourier Temporal Pyramid for robust spatio-temporal representation, enhancing the discriminative power of the model for action recognition. The proposed method demonstrates significant improvements over existing methods, achieving higher accuracies on both multiview and single-view benchmark datasets, thereby marking a substantial advancement in the field of 3D action recognition.
Xin et al. [67] introduces a novel approach to action recognition in video, addressing the complexities of real-world environments. The core of the paper is the development of the Adaptive Recurrent-Convolutional Hybrid (ARCH) Networks, a hybrid deep neural network (DNN) that integrates static, dynamic, and sequential mixed features for robust action recognition. This approach is designed to tackle three key challenges: spatial domain variations (such as noise and occlusion), temporal domain polytropes (differences in action velocities and durations), and intra- and inter-class diversities (similarities and differences within and between action categories). The ARCH model, evaluated on the UCF101 dataset and further validated through transfer learning on other datasets, demonstrates superior performance in recognizing human actions, outperforming existing deep-learning-based methods. This advancement is significant in its comprehensive handling of the spatial, temporal, and class diversity challenges in action recognition, offering a more effective and adaptable solution for long-term action recognition in uncontrolled and realistic environments.
By 2017, CNNs had become a key technology in the field of action recognition within computer vision. Their strength lay in their ability to autonomously extract meaningful spatial features from images or video frames. However, the complexity of action recognition required not just spatial but also temporal analysis. This led to the integration of CNNs with architectures capable of temporal processing, like RNNs or LSTMs, or the development of specialized CNN variants like 3D CNNs. Challenges such as handling diverse and complex video data, as well as the computational demands of training these models, were significant. Nonetheless, the advancements in CNN-based action recognition systems opened up numerous practical applications, ranging from surveillance to interactive gaming, and continued to be a vibrant area of research and development in the field of computer vision.
Zhao et al. [68] introduces a novel approach for action recognition in 3D videos, a field with significant applications in health monitoring, assisted living, and surveillance. The study’s core innovation is the integration of RNNs and CNNs in a unique two-stream fusion method. The RNNs, utilizing gated-recurrent-units, excel in processing long-term information from optical tracking data, while the 3D-CNNs are adept at handling detailed, recent information from video data. This combination allows for a comprehensive analysis of both spatial and temporal data aspects. The paper employs the NTU RGB+D Dataset, one of the largest datasets for 3D action recognition, containing over 56,000 action samples in 60 different classes. The novelty of the research lies in its two-stream fusion approach, which merges the outputs of RNN and CNN models, significantly enhancing action recognition performance. This approach includes two fusion structures: decision fusion and feature fusion, with the latter demonstrating superior results. The proposed method shows a remarkable improvement over state-of-the-art methods, outperforming them by more than 14% in both cross-subject and cross-view settings. Additionally, the novel RNN structure proposed in the study is optimized for faster convergence and reduced computational demand, making it a significant advancement in the field of action recognition in 3D videos.
Ullah et al. [69] presents an innovative approach to action recognition in video sequences, a challenging area in computer vision. The core of their methodology is the integration of Deep Bidirectional Long Short-Term Memory (DB-LSTM) networks with CNN features. This combination is designed to effectively capture and analyze the sequential information in video frames, addressing key challenges such as visual content similarity, changes in viewpoint, and varying illumination conditions. The model utilizes deep features extracted from every sixth frame using a pre-trained AlexNet, significantly reducing redundancy and computational complexity. The method is rigorously evaluated on three benchmark datasets: UCF-101, YouTube 11 Actions, and HMDB51, demonstrating its effectiveness in recognizing a wide range of actions. The novelty of this research lies in its efficient use of DB-LSTM for processing sequential frame changes in conjunction with CNN for deep visual data analysis, making it particularly adept at handling lengthy videos. The experimental results show that this approach significantly outperforms existing state-of-the-art methods in action recognition, highlighting its potential for applications in areas like surveillance, human-computer interaction, and sports analysis.
Li et al. [70] introduced an innovative approach for action recognition in 3D skeleton videos. The core of this research is the development of a method that transforms 3D skeleton data into color images, ensuring translation-scale invariance and independence from the dataset. This transformation is crucial as it effectively handles variations in translation and scale, common challenges in 3D skeleton data analysis. The study then employs a multi-scale deep CNN, built upon pre-trained models like ResNet, to classify these transformed images. This multi-scale CNN is specifically designed to enhance the model’s ability to adjust temporal frequencies, a key aspect when dealing with action recognition. The methodology is rigorously tested on the NTU RGB-D dataset, known for its complexity and diversity, and the results demonstrate that the proposed approach significantly outperforms existing methods. The novelty of this work lies in its unique image mapping technique and the application of a multi-scale deep CNN, making it a substantial contribution to the field of action recognition from 3D skeleton videos.
Varol et al. [71] introduced a novel approach to action recognition in videos through the use of Long-term Temporal Convolutions (LTC) within CNNs. Addressing the limitation of existing methods that typically analyze actions over short video intervals, this study proposes a model that captures the full temporal extent of human actions, which often last several seconds and encompass a wide range of frames. The LTC-CNN model comprises five space-time convolutional layers followed by three fully connected layers, designed to extend the temporal reach of the network while balancing complexity by reducing spatial resolution. This approach is evaluated on two challenging HAR benchmarks, UCF101 and HMDB51, known for their diverse and complex action categories. The paper’s innovation lies in its focus on long-term temporal patterns, a critical aspect often overlooked in action recognition. The study not only demonstrates the effectiveness of LTC in capturing extensive temporal information but also underscores the importance of high-quality optical flow estimation for accurate video representation. The results show state-of-the-art performance on both datasets, highlighting the model’s superiority in handling the temporal dynamics of human actions.
In 2018, Zhigang Tu et al. [72] introduces a novel approach to video-based HAR using a multi-stream CNN architecture. This model, termed Human-Related Multi-Stream CNN (HR-MSCNN), innovatively combines three motion-related and three appearance-related streams to analyze human actions in videos. The model employs Block-Sparse Robust Principal Component Analysis (B-RPCA) for effective human detection, addressing challenges like background motion and illumination changes, and uses a motion saliency measure to refine foreground regions and identify primary moving body parts. Evaluated on benchmark datasets such as JHMDB, HMDB51, UCF Sports, and UCF101, the HR-MSCNN demonstrates state-of-the-art performance. The paper’s novelty lies in its end-to-end CNN that considers both the detected person and discriminative motion regions, a new method for obtaining static and dynamic features, and a fusion method to efficiently combine information from six streams for action prediction. The findings highlight the model’s enhanced ability to focus on the most informative parts of the human body for accurate action recognition, marking a significant advancement in the field.
Arunnehru et al. [73] introduced an approach to improving surveillance systems through the application of 3D-CNNs integrated with 3D motion cuboids. This method is designed to enhance the accuracy of suspicious action recognition in video surveillance, a critical aspect of ensuring public safety and preventing potential crimes. By utilizing the spatial-temporal data encapsulated in 3D motion cuboids and processing this information through 3D-CNNs, the model is capable of detecting and recognizing complex human actions with higher precision than previously established methods. The research is validated through extensive testing on two benchmark datasets, KTH and Weizmann, where it demonstrated superior performance in terms of accuracy, thereby outperforming existing action recognition techniques. This advancement not only marks a significant contribution to the fields of intelligent video surveillance and computer vision but also opens new avenues for the development of more sophisticated and reliable surveillance systems capable of preemptively identifying and mitigating risks associated with suspicious activities.
Miao Ma et al. [74] introduced a novel approach for recognizing both general and fine-grained human actions in video sequences using a six-stream CNN. This method significantly focuses on fine-grained actions, which are more subtle and occur in smaller image regions compared to general actions. The core of the approach involves estimating human poses and body parts positions in video sequences, followed by cropping image patches at various scales to capture diverse appearance and motion cues. Uniquely, the model employs the outputs of the pooling layer of the CNN, which retains more spatial information than the fully connected layers typically used in such analyses. This is complemented by an effective encoding method that generates discriminative action descriptors for classification. The model’s efficacy is demonstrated on two public datasets, the Sub-JHMDB and MPII Cooking datasets, where it outperforms several state-of-the-art methods. The novelty of this work lies in its combination of coarse human pose tracking with multi-scale patch extraction and advanced pooling layer encoding, making it a significant contribution to the field of fine-grained action recognition in videos.
Zhang et al. [75] introduces a novel approach to real-time action recognition in videos. Addressing the limitations of traditional two-stream CNNs, which are hindered by the time-consuming computation of optical flows (OFs), this study proposes the use of motion vectors (MVs) as a more efficient alternative. MVs, inherent in video streams, can be extracted with minimal computational overhead, offering a pathway to real-time processing. However, MVs inherently contain noise and lack the fine details present in OFs, leading to potential accuracy degradation when used directly in CNNs. To overcome this, the paper presents four innovative training strategies that leverage the knowledge from OF-based CNNs to enhance the performance of MV-based CNNs. This knowledge transfer is rooted in the structural similarities between MV and OF, enabling the application of insights gained from OF CNNs to improve the generalization ability of MV CNNs. The core of this approach is the deeply transferred MV CNN, which employs a method called ”Deeply Connected Transfer” for multi-layer knowledge transfer from OF CNN to MV CNN. This method significantly boosts the accuracy of MV CNNs while maintaining the high-speed processing capabilities essential for real-time action recognition. The effectiveness of this approach is demonstrated through extensive experiments on various datasets, including UCF101, HMDB51, and THUMOS14, showcasing the model’s ability to achieve high accuracy and processing speed, thus marking a significant advancement in the field of real-time video-based action recognition.
Liu et al. [76] introduced a novel architecture for efficient and real-time action recognition in videos. The core of this architecture is the Temporal Convolutional 3D Network (T-C3D), which innovatively combines a residual 3D CNN with a temporal encoding method to capture both the appearance of individual frames and the motion between consecutive frames in a video. This approach addresses the computational inefficiencies of existing methods, particularly those reliant on optical flow extraction, by focusing on hierarchical multi-granularity learning. The T-C3D model is evaluated on two widely recognized action recognition datasets, UCF101 and HMDB51, where it demonstrates significant improvements in both accuracy and processing speed compared to state-of-the-art real-time methods. Notably, the T-C3D achieves over 5.4% higher accuracy and is twice as fast in terms of inference speed, reaching up to 969 frames per second. The model’s ability to process RGB frames in real-time while maintaining high accuracy and capturing both short-term and long-term temporal dynamics marks a substantial advancement in the field of video-based action recognition. The paper’s contribution lies in its novel integration of 3D-CNN with temporal encoding, enabling efficient modeling of video dynamics and facilitating real-time application potential.
In 2019, Jiang et al. [77] an innovative approach for video action recognition by efficiently integrating spatiotemporal and motion features within a unified 2D CNN framework, a significant departure from the traditional reliance on 3D convolution and optical flow pre-calculation. Central to this approach is the STM network, which comprises two novel components: the Channel-wise SpatioTemporal Module (CSTM) and the Channel-wise Motion Module (CMM). The CSTM focuses on capturing spatiotemporal features through channel-wise 1D convolution, allowing for effective temporal information fusion with minimal computational overhead. In contrast, the CMM is designed to encode motion features by learning distinct motion characteristics for each channel between adjacent time steps. This method was rigorously tested on various datasets, including Something-Something v1 & v2, Jester, Kinetics-400, UCF-101, and HMDB-51, demonstrating superior performance over existing state-of-the-art methods. The novelty of this research lies in its unique integration of spatiotemporal and motion features in a 2D framework, efficiently handling different types of action recognition scenarios and showcasing its potential in practical applications.
Crasto et al. [78] introduced a groundbreaking approach to action recognition by ingeniously integrating motion information into a standard 3D CNN that processes RGB frames, thereby circumventing the computationally intensive and latency-prone process of calculating optical flow, a common necessity in conventional two-stream action recognition frameworks. Central to this innovation are two novel learning strategies: MERS (Motion-Emulated RGB Stream) and MARS (Motion-Augmented RGB Stream). MERS trains the network to emulate features from the motion stream, significantly reducing the computational burden, while MARS further advances this concept by synergistically combining appearance and motion information into a singular, efficient stream. This approach not only achieves superior performance, outstripping individual RGB and Flow streams with 72.7% accuracy on the Kinetics dataset and 80.1% on HMDB51 (split-1), but also maintains the high standards of two-stream architectures without their inherent computational inefficiencies. The paper’s innovative use of a feature-based loss, alongside the standard cross-entropy loss, for network training marks a significant leap forward, showcasing the effective integration of temporal information into a 3D CNN and setting a new benchmark for future research in action recognition.
Hussein et al.[79] introduced a novel framework aimed at enhancing the recognition of complex human activities in videos. Recognizing the limitations of traditional spatiotemporal 3D convolutions, which struggle with the variability and length of complex actions, the authors propose Timeception convolution layers. These layers are uniquely designed to handle minute-long temporal patterns, extending the temporal modeling capacity up to 1024 timesteps, significantly surpassing existing methods. Timeception achieves this by employing multi-scale temporal convolutions and reducing the complexity inherent in 3D convolutions, making it particularly adept at capturing the nuances of complex actions that consist of multiple, variably-timed one-actions. The model’s efficacy is demonstrated through impressive accuracy on challenging datasets like Charades, Breakfast Actions, and MultiTHUMOS, marking a substantial advancement in the field of action recognition. The paper’s innovative approach in decomposing and efficiently managing the temporal aspect of videos paves the way for more nuanced and accurate video analysis, particularly in understanding and interpreting complex, multi-faceted human activities.
He et al. [80] presents a sophisticated approach to action recognition in videos through the novel StNet architecture, which adeptly captures both local and global spatial-temporal relationships. StNet ingeniously stacks successive video frames into a super-image and applies 2D convolution to efficiently encode local spatial-temporal information, a method that contrasts with traditional CNN+RNN or pure 3D convolution-based approaches. To grasp the global spatial-temporal structure, the architecture incorporates temporal convolutions and introduces the innovative Temporal Xception Block (TXB), which separately processes channel-wise and temporal-wise convolutions, optimizing the model for size and computational efficiency. Evaluated on the Kinetics dataset, StNet demonstrates superior performance over existing 2D and 3D convolution-based solutions, striking a commendable balance between recognition accuracy and model complexity. Furthermore, the model’s robustness and generalizability are underscored by its successful application to the UCF101 dataset, showcasing its potential in advancing the field of video-based action recognition.
Hao Yang et al. [81] presented a novel approach to action recognition, addressing the computational and storage challenges associated with traditional 3D CNNs. The authors introduce an innovative model that utilizes efficient asymmetric one-directional 3D convolutions to approximate the conventional 3D convolutions, significantly reducing the number of parameters and computational cost. Central to the model are MicroNets, a series of local 3D convolutional networks incorporating multi-scale 3D convolution branches, designed to enhance the feature learning capacity of asymmetric 3D convolutions without increasing computational demands. The model also features a multi-source enhanced input method, which fuses information from RGB and Flow frames at the pre-processing stage, thereby circumventing the need for training separate networks for each frame type and further reducing computational requirements. Evaluated on two challenging benchmarks, UCF-101 and HMDB-51, the asymmetric 3D-CNN model not only outperforms traditional 3D-CNN models in terms of efficiency and effectiveness but also achieves comparable results with state-of-the-art action recognition methods, marking a significant advancement in the field of computer vision.
In 2020, Sudhakaran et al. [82] introduces a novel approach to video action recognition by proposing the Gate-Shift Module (GSM), a lightweight and efficient module designed to transform standard 2D CNNs into powerful spatio-temporal feature extractors. The GSM innovatively applies spatial gating to the spatial-temporal decomposition of 3D kernels, allowing the network to adaptively route and combine features through time with minimal additional computational cost and parameter overhead. This approach addresses the challenges of the high dimensionality and computational complexity typically associated with 3D CNNs in video action recognition. The paper demonstrates the effectiveness of the GSM through extensive evaluations, where the model achieves state-of-the-art results on the Something Something-V1 and Diving48 datasets and competitive results on the EPIC-Kitchens dataset, all while maintaining a significantly lower model complexity compared to existing methods. This breakthrough presents a significant advancement in the field, offering a more efficient and scalable solution for video action recognition tasks.
Liu x et al. [83] investigated the application of the VGG-16 Convolutional Neural Network model to the recognition of human actions from video data. This research employs deep learning techniques to analyze kinematic behavior, utilizing two datasets: the KTH dataset, featuring actions like walking and boxing, and the Weizmann dataset, which includes activities such as bending and waving. The study’s novelty lies in adapting the VGG-16 model, typically used for static image recognition, to dynamically recognize human behaviors, highlighting its efficacy compared to traditional methods like the histogram of optical flow orientation. Findings indicate that the VGG-16 model significantly outperforms other methods, achieving an average accuracy of 91.93% on the KTH dataset and 95.75% on the Weizmann dataset, with particularly high accuracy in recognizing boxing and bending movements. These results underscore the potential of advanced CNN models in the real-time analysis of complex human movements, paving the way for improvements in fields such as sports analytics and health monitoring.
Xiong et al. [84] presented a novel approach to enhance Human-Robot Collaboration (HRC) in smart manufacturing by improving the accuracy and reliability of HAR. Recognizing the challenges in effectively leveraging temporal information and the scarcity of large training datasets in manufacturing environments, the paper introduces an integrated method that combines optical flow and CNN-based transfer learning. Optical flow images, encapsulating the temporal dynamics of human motion, are fed into a two-stream CNN structure, meticulously designed to parse both spatial and temporal information concurrently. To circumvent the issue of limited training data, the study explores transfer learning, enabling the adaptation of a pretrained CNN to the specific context of manufacturing scenarios. The efficacy of this innovative approach is validated through its application to engine block assembly tasks, showcasing its potential in advancing the capabilities of HRC systems and contributing significantly to the evolution of smart manufacturing practices. The paper’s novel integration of optical flow with a two-stream CNN, augmented by transfer learning, marks a substantial stride in addressing the intricate demands of HAR within the realm of industrial automation.
Sahoo et al. [85] innovates in HAR by merging depth information with sequential learning. The HAR-Depth model integrates a Deep Bidirectional Long Short Term Memory (DBiLSTM) for capturing the temporal dynamics of actions and employs depth history images (DHI) to enhance shape recognition, where depth-estimated frames are projected onto the X-Y plane. Importantly, the model harnesses CNNs for initial feature extraction from action frames, which are then processed through the DBiLSTM. Evaluated on datasets such as KTH, UCF Sports, JHMDB, UCF101, and HMDB51, the model demonstrated superior accuracy, notably achieving 97.67% on KTH and 95.00% on UCF Sports. This approach highlights the model’s capability to distinguish similar actions through an effective combination of depth-based shape information and sophisticated temporal sequence processing, positioning it as a potent solution for advanced surveillance and interactive systems.
[…]

### v2

@@ 3.1.1 Two-Stream CNNs
[81] presented an innovative approach to recognize actions in video sequences by using a two-stream CNN architecture. This approach divides the task into two distinct problems: recognizing spatial features from single frames and capturing temporal features across frames. The spatial stream CNN processes static visual information, while the temporal stream CNN handles motion by analyzing optical flow. The model was tested on benchmark datasets like UCF-101 and HMDB-51, where it achieved state-of-the-art results, showcasing the effectiveness of this two-stream method. The novelty of this work lies in the separation of motion and appearance features, which allows for more specialized networks that can better capture the complexities of video-based action recognition. The success of this model has made a significant impact on the field, influencing many future research directions in video understanding. Consequently, numerous methods have been proposed to enhance the the two-stream model [99, 33, 100, 75, 104].
In 2016, building on the the two-stream CNN, [33] focused on improving the two-stream CNN by exploring various fusion strategies for combining spatial and temporal streams, resulting in better performance on the UCF-101 and HMDB-51 datasets. By enhancing fusion techniques, this work addressed the limitations of the initial two-stream model, leading to more effective integration of spatial and temporal information. [100] introduced temporal segment networks (TSN). This work aimed to capture long-range temporal structures for action recognition, achieving significant improvements on the UCF-101 and HMDB-51 datasets by dividing videos into segments for comprehensive analysis. The introduction of TSN extended the temporal analysis capabilities of the two-stream CNN, enabling the capture of long-range dependencies.
In 2017, derived from the two-stream CNN, [20] proposed a three-stream method by using spatio-temporal vectors, with locally max-pooled features to enhance performance. Tested on the UCF-101 and HMDB-51 datasets, the approach demonstrated improved recognition accuracy by efficiently capturing spatio-temporal dynamics. In 2018, the efficient convolutional network for online video understanding (ECO) was introduced by [114], combining the two-stream CNN approach with lightweight 3D CNNs, and focusing on efficiency and real-time processing, with high efficiency and competitive accuracy demonstrated on the Kinetics and UCF-101 datasets.
[32] introduced the SlowFast network which processes video data at varying frame rates to capture both spatial semantics and motion dynamics, achieving state-of-the-art results on the Kinetics-400 and Charades datasets. By introducing different temporal resolutions, this work innovated on the two-stream concept, capturing fine and coarse temporal details. [101] expanded on their previous work with TSN, developing a multi-stream approach that incorporated RGB, optical flow, and warped optical flow streams to model long-range temporal structures more effectively. This approach achieved state-of-the-art results by capturing both spatial and temporal information across various time scales. In 2021, temporal difference networks (TDN) were introduced by [98], leveraging the multi-stream CNN with a focus on capturing motion dynamics efficiently. Using the UCF-101 and HMDB-51 datasets, TDN achieved notable improvements by effectively modeling temporal differences. By emphasizing temporal differences, this work advanced the ability of the two-stream CNN to capture motion dynamics more effectively.
Table LABEL:table:TwoStreamNetworks presents the works discussed in this section that utilized two or more stream CNNs approaches.
Table 1: Two-stream CNN-based approaches in HAR.
Paper Model Dataset Novelty
[81] Two-stream CNN UCF-101, HMDB-51 Introduced the two-stream architecture separating spatial and temporal streams for effective action recognition.
[33] Two-stream CNN UCF-101, HMDB-51 Explored various fusion strategies to combine spatial and temporal streams, and improved performance.
[100] Two-stream CNN + TSN UCF-101, HMDB-51 Introduced TSN to capture long-range temporal structures by dividing videos into segments.
[20] Three-Stream CNN UCF-101, HMDB-51 Proposed a three-stream method using spatio-temporal vectors with locally max-pooled features for enhanced performance.
[114] Two-stream CNN + 3D CNN Kinetics, UCF-101 Combined the two-stream CNN with lightweight 3D CNNs for efficient real-time processing.
[32] Two-stream CNN + SlowFast Kinetics-400, Charades Introduced SlowFast networks processing video data at varying frame rates to capture both spatial and motion dynamics.
[101] CNN-RNN, (Multi-stream TSN) UCF101, HMDB51 Expanded on TSN by developing a multi-stream approach that incorporated RGB, optical flow, and warped optical flow streams to model long-range temporal structures more effectively.
[98] Multi-stream CNN + TDN Something-Something V1 and V2 Introduced TDN focusing on capturing motion dynamics efficiently.
3.1.2 3D CNN-Based Approaches
The foundational work conducted by [49] introduced 3D CNNs for HAR, demonstrating their effectiveness in capturing spatio-temporal features on the KTH and UCF-101 datasets and outperforming traditional 2D CNNs. The work paved the way for further research on enhancing 3D convolutional models. [90] introduced C3D, a generic 3D CNN for spatio-temporal feature learning, achieving state-of-the-art performance on the Sports-1M and UCF-101 datasets and highlighting the scalability and effectiveness of 3D convolutions. Building on the work by [49], C3D demonstrated the potential of 3D CNNs across diverse datasets, influencing subsequent research in 3D CNNs. [95] introduced long-term temporal convolutions to capture extended motion patterns. This work improved the accuracy on the UCF-101 and HMDB-51 datasets and emphasized the importance of long-term motion information. Moreover, this study extended the temporal scope of 3D CNNs, highlighting the need for capturing long-term motion for accurate action recognition. In the same year, [76] proposed pseudo-3D residual networks (P3D), which combined 2D and 3D convolutions to balance the accuracy and computational complexity. This work achieved competitive performance on the Kinetics and UCF-101 datasets. Moreover, P3D networks offered a more efficient approach by blending 2D and 3D convolutions, further refining the capabilities of 3D CNNs. Additionally, [15] introduced I3D by inflating 2D convolutions to 3D, achieving significant improvements on the Kinetics dataset by leveraging ImageNet pre-training, thereby setting new performance benchmarks. I3D bridged the gap between 2D and 3D CNNs, demonstrating the benefits of transfer learning in 3D convolutional models.
[42] evaluated the scalability of 3D CNNs with increased data and model sizes, demonstrating that deeper 3D CNNs can achieve better performance on the Kinetics and UCF-101 datasets, paralleling the success of 2D CNNs on ImageNet. This study emphasized the need for larger datasets and deeper models in 3D convolutional research, highlighting the potential of 3D CNNs to retrace the historical success of 2D CNNs. Building on these insights, [26] introduced a new temporal 3D ConvNet architecture with enhanced transfer learning capabilities, demonstrating superior performance on the UCF-101 and HMDB-51 datasets through architectural innovations and effective transfer learning. This work underscored the importance of architectural innovation and transfer learning, pushing the boundaries of 3D CNN performance and further advancing the field of action recognition. [92] further contributed by conducting a comprehensive analysis of spatio-temporal convolutions, highlighting the benefits of factorizing 3D convolutions into separate spatial and temporal components, achieving state-of-the-art results on the Kinetics and UCF-101 datasets. This dissection provided insights that informed subsequent model designs and optimizations. In the same year, [107] explored the trade-offs between speed and accuracy in spatio-temporal feature learning, proposing efficient 3D CNN variants that balance computational cost and recognition performance on the Kinetics and UCF-101 datasets. Their work highlighted the practical considerations of deploying 3D CNNs, emphasizing the need to balance speed and accuracy, thereby refining the approach to spatio-temporal feature learning. Additionally, [103] introduced non-local neural networks to capture long-range dependencies, demonstrating that non-local operations significantly improve the modeling of complex temporal relationships and enhance action recognition performance on the Kinetics and Something-Something datasets. By integrating non-local operations, this study advanced the ability of 3D CNNs to capture complex temporal patterns, further pushing the boundaries of spatio-temporal modeling.
[…]